### v1

# Skyrmion motion in a synthetic antiferromagnet driven by asymmetric spin wave emission

Abstract
Skyrmions have been proposed as new information carriers in racetrack memory devices. To realise such devices, a small size; high speed of propagation; and minimal skyrmion Hall angle are required. Synthetic antiferromagnets (SAFs) present the ideal materials system to realise these aims. In this work, we use micromagnetic simulations to propose a new method for manipulating them using exclusively global magnetic fields. An out-of-plane microwave field induces oscillations in the skyrmions radius which in turn emits spin waves. When a static in-plane field is added, this breaks the symmetry of the skyrmions and causes asymmetric spin wave emission. This in turn drives motion of the skyrmions, with the fastest velocities observed at the frequency of the intrinsic out-of-phase breathing mode of the pair of skyrmions. This behaviour is investigated over a range of experimentally realistic antiferromagnetic interlayer exchange coupling strengths, and the results compared to previous works studying similar motion driven with an oscillating electric field. Through this the true effect of varying the exchange coupling strength is determined, and greater insight is gained into the mechanism of skyrmion motion. These results will help to inform the design of future novel computing architectures based on the dynamics of skyrmions in synthetic antiferromagnets.
I Introduction
Magnetic skyrmions—a type of topological spin texture stabilised by the Dzyaloshinksii-Moriya interaction [1, 2, 3]—are an area of great recent interest for their potential applications in magnetic memory [4, 5], future computation [6, 7] and quantum operations [8, 9]. In all of these applications, the ability to controllably move and manipulate skyrmions is key. They can be moved with electrical currents [10, 11, 12], magnetic field gradients [13, 14], and microwave magnetic fields [15, 16] among other methods. However, in ferromagnetic systems, skyrmions exhibit the skyrmion Hall effect—where they travel at an angle to the driving force [17]—the magnitude of which depends on the velocity, skyrmion size and local pinning landscape [18, 19, 12, 20]. This is a significant roadblock to the implementation of skyrmion based devices as skyrmions will tend to drift to the edges of devices and annihilate [21].
Synthetic antiferromagnets—where two individually ferromagnetic layers are sandwiched around a nonmagnetic spacer layer and coupled antiferromagnetically [22]—have been predicted to be an advantageous platform for skyrmions [23, 21]. This is because of their proposed smaller size, faster velocity, and increased resistance to stray fields compared to skyrmions in ferromagnetic systems [24]. Skyrmions with radii as small as $\sim$50 nm have been experimentally in synthetic antiferromagnets [25, 26]. Skyrmions and domain walls have been shown to move at substantially faster velocities than their ferromagnetic equivalents of up to 900 m/s [27] in the case of skyrmions, or 750 m/s in the case of domain walls [28]. Domain walls have also been shown to depin at current densities as low as $\mathrm{1\times 10^{11}\ A/m^{2}}$ [29, 30]. Due to the challenges in detection, much of the body of research on skyrmions in synthetic antiferromagnets remains theoretical [31], and experimental measurements of skyrmion motion are extremely limited [26, 27, 32].
Skyrmions have been shown to exhibit microwave resonances when excited by magnetic and electric fields [33]. This was originally predicted based on numerical modelling of skyrmions in bulk systems [34], and was first demonstrated experimentally in Cu2OSeO3 [35]. The breathing modes of skyrmions were first simulated in thin films by Kim et al. [36], and have recently been demonstrated experimentally in ferromagnetic thin films [37, 38]. In synthetic antiferromagnets, micromagnetic simulations have been used to demonstrate both the breathing modes [39, 40] and corresponding rotational modes [41]. By exciting these breathing modes using either electric, or magnetic fields, it has been demonstrated that skyrmions in ferromagnetic systems will emit spin waves asymmetrically in a static in-plane field, and this stimulates skyrmion motion with a peak in velocity as a function of frequency that is coupled to the peak of the breathing mode amplitude [16, 42]. This has also been demonstrated in synthetic antiferromagnets where microwave electric fields were used to excite straight-line skyrmion motion [43]. In that work, a dependence of the velocity on the strength of the antiferromagnetic interlayer exchange coupling was demonstrated, however the link between this and the skyrmion breathing modes was not discussed [43]. This microwave field driven approach to skyrmion motion has the advantage that—unlike the conventional method of using an electric current—it does not require a conductive sample, and so can be used to manipulate skyrmions in insulating materials.
In this work, we use microwave magnetic fields in micromagnetic simulations to drive skyrmion breathing modes in a synthetic antiferromagnet, and combine these with an in-plane, static field to excite skyrmion motion in a wiggle-like fashion. We investigate this motion as a function of frequency, and relate it to the intrinsic breathing modes of these skyrmions. We use this understanding of the coupling between velocity and breathing mode frequency to understand the true relationship between the skyrmion velocity and the antiferromagnetic interlayer exchange coupling strength. Finally, we use the additional knowledge of this mechanism to discuss the differences between the previous electric-field driven skyrmion motion [43] and our work. The additional understanding of the relationship between the skyrmion breathing mode and the velocity can be used in applications to modulate the speed of the skyrmion using only the microwave frequency. The emission of spin waves by a domain wall driven by microwave magnetic fields has also very recently been demonstrated [44]. This, together with the similarly recent imaging of spin waves in a synthetic antiferromagnet [45], paves the way to the experimental realisation of our method.
II Methods
The synthetic antiferromagnetic multilayer presented in this paper was studied using the micromagnetic solver MUMAX3 [46]. MUMAX3 is a finite difference numerical solver for the Landau-Lifshitz-Gilbert (LLG) equation that calculates the spatial and time dependence of the magnetisation $\mathbf{m}$ in a simulation [46]. The LLG equation is given by
$\dfrac{d\mathbf{m}}{dt}=-\lvert\gamma_{0}\rvert\mathbf{H}_{\mathrm{eff}}+\alpha\mathbf{m}\times\dfrac{d\mathbf{m}}{dt},$ (1)
where $\mathbf{m}$ is the magnetisation unit vector, $\alpha$ is the Gilbert damping parameter, and $\mathbf{H}_{\mathrm{eff}}$ is the effective magnetic field. The effective magnetic field is proportional to the differential of the total micromagnetic energy $U$, given by
$\mathbf{H}_{\mathrm{eff}}=-\dfrac{1}{\mu_{0}M_{\mathrm{s}}}\dfrac{\delta U}{\delta\mathbf{m}}.$ (2)
Here $\mu_{0}$ is the permeability of free space which is $4\pi\times 10^{-7}\ \mathrm{N\cdot A^{-2}}$, and $M_{\mathrm{s}}$ is the saturation magnetisation of the system. The micromagnetic energy $U$ includes the usual energy terms that describe the interaction with an external magnetic field, the effects of the demagnetisation field, the Heisenberg exchange. In these simulations we consider a single anisotropy energy term perpendicular to the plane of the field. This is given by
[…]
where $K_{\mathrm{u}}$ is the anisotropy constant, and $\mathbf{u}$ is a unit vector denoting the direction of the ansiotropy, in this case along the vector perpendicular to the film plane. In addition we consider an interfacial DMI term of the form
[…]
where D is the DMI constant. Finally, we add a custom energy term to the simulation to model the indirect interlayer exchange coupling, given by
[…]
where the subscript of $\mathbf{m}$ denotes the layer number, and $\sigma_{\mathrm{RKKY}}$ is the coupling energy. The system is antiferromagnetic for $\sigma_{\mathrm{RKKY}}<0$.
The simulation is set up with exchange stiffness $A=15\times 10^{-12}$ J/m, saturation magnetisation $M_{\mathrm{s}}=580\times 10^{3}$ A/m, uniaxial anisotropy constant $K=7\times 10^{5}\ \mathrm{J/m^{3}}$, interfacial DMI $D=3\times 10^{-3}\ \mathrm{J/m^{2}}$ and Gilbert damping parameter $\alpha=0.01$. These values represent typical generic parameters of metallic multilayers with DMI that are usually chosen for synthetic antiferromagnets hosting skyrmions [47]. They are also chosen specifically to mirror those used in ref. [43] in order to better compare results. The system was initialised with an in plane x-y cellsize of 1.5625 nm and an out-of-plane z cellsize of 0.4 nm. The configuration of cells was $256\times 256\times 4$ cells, for a total layer width and length of 400 nm and a total stack thickness of 1.6 nm. Only two of the layers were magnetic, with the remaining two layers not contributing to the simulation other than to mirror the spatial separation of the magnetic layers as in a real system.
To measure the frequency response of the system, it was excited at $t=0$ with a sinc field along the z-axis with a cutoff frequency of $f$ = 100 GHz of the form
$H_{z}=H_{0}+H_{1}\dfrac{\mathrm{sin}(2\pi ft)}{2\pi ft},$ (6)
where $H_{0}$ is an optional static field, $H_{1}$ is the magnitude of the sinc field, 0.5 mT unless otherwise specified. This was used as it is a top-hat function in Fourier space, and so excites all modes equally up to its cutoff frequency. Simulations were then run for 10-20 ns, and the magnetisations of each layer were recorded. The change in magnetisation, based on the spatial average of the magnetisation in each layer at time $t$ $\langle m_{z}(t)\rangle$ is calculated as
$\delta m_{z}=\langle m_{z}(t)\rangle-\langle m_{z}(0)\rangle.$ (7)
The frequency response of the system is then calculated using the power spectral density (PSD), which is given by the Fourier transform of $\delta m_{z}$ multiplied by its complex conjugate
[…]
To excite the skyrmion motion—unless otherwise specified—a static in-plane field of 0.5 T is used, together with a static out-of-plane field of 50 mT and an out-of-plane microwave excitation field of 20 mT. The microwave excitation field is smaller than that used by other works [15] but a comparable order of magnitude. In order to track the motion of the skyrmion, for each timestep we calculate the guiding centre $\mathbf{R}_{\mathrm{Sk}}$ of each magnetic layer, given by
[…]
where $\mathbf{r}$ is the position vector of each cell, and q is the topological charge density
$q=\mathbf{m}\cdot\Big{(}\dfrac{\partial\mathbf{m}}{\partial x}\times\dfrac{\partial\mathbf{m}}{\partial y}\Big{)},$ (10)
[…]
III Static and Dynamic Characterisation
III.1 Effect of in-plane field on skyrmions
Figure 1: Illustration of simulation set-up and characterisation of skyrmion radius as a function of field. (a) Illustration of the simulated multilayer. Two ferromagnetic (FM) layers, labelled FM1 and FM2 are coupled together antiferromagnetically through a non-magnetic (NM) spacer layer. An in-plane field ($H_{\mathrm{IP}}$) causes the skyrmion to elongate, while the sinusoidal driving field causes it to emit spin waves and move perpendicular to $H_{\mathrm{IP}}$. (b) and (c) show the cross sections of the two skyrmions in layers 1 and 2 along the x-axis in an in-plane field along the x-axis of 1.7 T. The effect of the in plane field is shown on the radius of the two coupled skyrmions for field along the y-axis in (d) and field along the x-axis in (e). The ratio of the radius along the x- and y-axes of each skyrmion for each field orientation is shown in (f). The quanization is related to the cell size of the simulations.
The effect of a static in-plane field on the skyrmion radius is shown in Fig. 1(b-e) for an antiferromagnetic interlayer exchange coupling strength of $-3\times 10^{-4}\ \mathrm{J/m^{2}}$ in a static out-of-plane field of 50 mT. For this antiferromagnetic interlayer exchange coupling strength, the interlayer coupling is sufficiently strong to overcome the effects of the out-of-plane field and so the two skyrmions have equal radius in zero in-plane field $H_{\mathrm{IP}}$. As the field increases above $\sim$0.3 T, the skyrmions in each layer begin to both stretch and shrink. This shrinking is caused by the in-plane field pulling the cores of the two skyrmions apart laterally—as shown in Fig. 1(b,c)—which introduces a breaking of the antiferromagnetic interlayer coupling between spins inside the skyrmion. Thus, the skyrmion size shrinks to minimize this. Because of this, the stretching of the skyrmion is better demonstrated by the ratio between $r_{x}$ and $r_{y}$, as shown in Fig. 1(f). Above fields of $\sim$0.3 T this behaviour starts to change, and steadily changes from 0.5 T onwards, indicating the stretching of the two skyrmions along the direction of the field. On this basis, unless otherwise specified, an in-plane field of 0.5 T is chosen for further simulations which causes a modest deformation of the skyrmion, and also allows for better intercomparison of results with ref. [43]. Simulations of the full hysteresis loops are shown in supplementary note LABEL:supp:note:hyst_loops. The saturation field along the x- and y-axes is 4.3 T, demonstrating that this in-plane field is only a modest $\sim$10 % of saturation.
III.2 Skyrmion breathing modes
Figure 2: Skyrmion frequency response as a function of frequency and antiferromagnetic interlayer exchange coupling strength. (a) Shows the frequency response for an interlayer exchange coupling strength of $-3\times 10^{-4}\ \mathrm{J/m^{2}}$. Panels (b-d) show colour maps of the response as a function of interlayer exchange coupling strength, denoted as $\mathrm{|\sigma_{RKKY}|}$. (b) Shows the response of the entire simulated system, whereas (c) and (d) show the response when calculating the PSD of just the magnetisation of the cells of layer 1 and layer 2, respectively. (b), (c) and (d) have a common colour bar.
Before the motion of the skyrmions is induced, we first study the breathing modes of the system following the approach of Ref. [39]. As outlined in the methods, the system is first initialised with a Néel-type skyrmion in each layer and allowed to relax in a static out-of-plane field of 50 mT. Then a sinc pulse of amplitude 0.5 mT and frequency 100 GHz is applied at time 0 and the simulation is run for 20 ns. The results of this are shown in Fig. 2. Fig. 2(a) shows an illustration of the skyrmion frequency response with an interlayer exchange coupling strength of $-0.3\ \mathrm{mJ/m^{2}}$ for the two individual layers as well as the whole system. Because periodic boundary conditions were used for these simulations, there is no curling of the magnetisation at the edge, and so there are no modes arising from this and the frequency response seen here comes entirely from the skyrmion [40]. We see at $\sim$14 GHz there is a mode present in each of the individual layers, whereas, in the total system, there is almost no response. At the higher frequency mode at $\sim$57 GHz, there is a response in all three plots. This has been well studied [39, 40]; the difference arises from the interplay between the two skyrmions at each mode. When excited at the lower mode, the two skyrmions breathe in phase with each other, and so at any time $t$, spins in each magnetic layer are coupled antiparallel over the whole lateral extent of the system, and so the net moment is always cancelled. Conversely at the higher frequency mode, the two skyrmions breathe in antiphase, and so at any time $t$ there is always a small net moment originating from the difference in size of these two skyrmions, and thus there is an observable peak in the frequency response of the system as a whole. Heatmaps of the frequency response of the whole system and the two individual layers are then shown in Fig. 2(b-d). There is a considerable variation in the frequencies of the two peaks, especially that of the second as a function of the interlayer antiferromagnetic exchange strength.
IV All magnetic field driven skyrmion motion
IV.1 Skyrmion motion driven by the addition of an in-plane magnetic field
Figure 3: In-plane field-driven skyrmion motion in a field of 0.5 T along x, driven at a frequency of 57.4 GHz. Panel (a) shows the x- and y-positions of the skyrmion guiding centre for both the skyrmions in layers 1 and 2 of the SAF. Panel (b) shows the calculated velocities of these guiding centre x- and y-positions as a function of time. Points on the lines in panels (a) and (b) are deliberately undersampled to see both the datapoints of layers 1 and 2. Panel (c) shows the trajectory of both skyrmions, with the y-positions of each skyrmion plotted against the x-position. The colour gradient of each line denotes the time, which is shown on the colour bar to the right of the panel. Arrow shows the direction of the in plane static field.
Considering the static and dynamic characterisation of skyrmions in the simulated system described in Sec. III.1 and Sec. III.2, we now move on to describing the effect of an in-plane field together with an excitation field on the motion of the SAF skyrmion. The frequencies of the two peaks of Fig. 2(a) are found by fitting the data to a Lorentzian lineshape to be 14.3 GHz and 57.4 GHz. A sinusoidal driving field is then used at the higher frequency to drive the skyrmion together with a static out-of-plane field of 50 mT as discussed in Sec. III.2. This is also done with an in-plane static field of 0.5 T. The results of this are shown in Fig. 3. Fig. 3(a) shows the x- and y-positions of the guiding centres of the skyrmions in each layer as a function of time for 20 ns. We see that—while the skyrmion moves steadily along the y-axis at a seemingly constant rate—the x-axis motion proceeds in an oscillatory fashion with an oscillation frequency of $\sim$1 GHz, with an additional very small net movement along the x-axis. This can also be visualised with the trajectories of the two skyrmions shown in Fig. 3(c). We see that the skyrmions move in a wiggle-like fashion, the first report of this kind of motion in a synthetic antiferromagnet. This is also unlike the motion observed in ref. [43] but similar to the trajectories in similar ferromagnetic systems reported in [42]. The position as a function of time is used to calculate the x- and y-velocities of each skyrmion as a function of time, which is shown in Fig. 3(b). The y velocities ramp up from 0 for the first $\sim$2 ns, before maintaining a steady value of $\sim$6 m/s for the rest of the simulation time. Meanwhile, the x velocities oscillate around 0, driving the wiggle-like motion. To calculate the average velocities presented later, an average is taken of the y-axis velocity once it reaches its steady value, here around 3 ns into the simulation time.
IV.2 Skyrmion velocity coupled to out-of-phase breathing
Figure 4: Velocities of the two skyrmions as a function of frequency of the driving field. Datapoints are deliberately undersampled so that both datasets are visible. The intensity of the breathing modes from Fig. 2 is overlaid with grey shading and rescaled to match the peak amplitudes. Inset shows the same schematic from Fig. 1 to illustrate the directions of the magnetic fields and the direction of skyrmion motion.
The average velocity as a function of frequency is presented in Fig. 4 for an antiferromagnetic interlayer exchange coupling strength of -0.3 $\mathrm{mJ/m^{2}}$, over a range of frequencies from 6 GHz to 97 GHz. Added for comparison is the PSD of the breathing modes from Fig. 2(a) rescaled to allow comparison with the calculated velocities. There is a strong overlap between the peak velocity of the skyrmion and the peak in the breathing mode intensity, suggesting a coupling between the breathing mode and the velocity. However, perhaps surprisingly there is no similar peak when driving the skyrmion motion at the lower frequency of the in-phase peak.
Between frequencies of $\sim$70 and $\sim$80 GHz in Fig. 4(a)—above the out-of-phase breathing resonant frequency—there is a second broader peak in velocity that is seemingly not related to any skyrmion breathing mode. This is reminiscent of the higher frequency hybridisation of skyrmion breathing with spin wave modes of ferromagnetic [36] and synthetic antiferromagnetic [39] nanodots that has been previously observed. However, as discussed in Sec. II, in this simulation periodic boundary conditions are used. This excludes such spin wave resonant modes, since they originate from spin canting at the edges of the simulation, which the application of periodic boundary conditions removes. We instead suggest that this higher frequency broad peak in velocity arises from exciting a hybrid mode between the skyrmion breathing and a spin wave mode arising from the canting of spins in the applied in-plane field. Because of the motion that would be induced by performing the simulations used to generate Fig. 2 in an in-plane field, this is difficult to show directly. However, in supplementary note LABEL:supp:sec:high_freq_mode we show and discuss simulation of the PSD in the presence of an in-plane field but without a skyrmion, which yields a peak of similar centre frequency and width to this second peak in velocity.
IV.3 The effect of the spin wave emission amplitude
Figure 5: (a,b) Snapshots of the $y$-component of the magnetisation in layers 1 and 2 respectively at t = 0.2 ns when driven at 57.4 GHz. The colormap has been windowed to show the spin wave emission. (c,d) Show line cuts through the centre of the skyrmion from (a,b), while (e,f) show the equivalent line cuts through the centre of the skyrmion in the case where the in-plane field is removed, but still at 0.2 ns after the start of the simulation with the same out-of-plane driving field. Panels (g-l) are all the same but with a 14.3 GHz driving field. (g,h) Are snapshots of layers 1 and 2 respectively at 0.2 ns, (i,j) are line cuts through (g,h), and (k,l) are the equivalent line cuts in the absence of an in-plane field. Panel (m) shows the average skyrmion velocity when driven with a static in-plane field of 0.5 T as a function of the out-of-plane excitation field. The gray shaded region denotes the field range below which there is no motion. Datapoints are deliberately undersampled so that both datasets are visible. Panel (n) shows the average peak-to-peak variation in the skyrmion radius when breathing as a function of the amplitude of the out-of-plane field in the absence of an in-plane field. The gray shaded region is taken from panel (m) and the black dashed line is a guide to the eye to separate radii measured with excitation fields above and below this region. Note that the error bars on these series are too small to be visible.
It has been discussed that the motion of the skyrmions in these systems is driven by spin wave emission [42, 16, 43], however, the mechanism of this spin wave emission has not been investigated in detail. This emission is represented more clearly in Figs. 5(a,b) which show the skyrmion when driven at 57.4 GHz in a 0.5 T in-plane magnetic field at 0.2 ns after the start of the simulation; at the onset of skyrmion motion. There is some interference present because of the periodic boundary conditions. Similarly Figs. 5(g,h) show the same situation when the system is driven with a field of 14.3 GHz, the frequency of the in-phase breathing mode. These images are complemented by supplementary videos S1 and S2 which show this spin wave emission over the first 1 ns of the simulation in the presence of an in-plane magnetic field. To further illustrate the spin-wave emission, each of these pairs of colormaps are complemented with line cuts that are take through the centre of the skyrmion, shown in Fig. 5(c,d) for the 57.4 GHz case, and Fig. 5(i,j) for the 14.3 GHz case. These line cuts are in turn complemented by supplementary videos S3 and S4, which again show this profile over the first 1 ns of the simulation, in an in-plane magnetic field of 0.5 mT, driven by an out-of-plane oscillating field of 20 mT, at 57.4 GHz and 14.3 GHz in supplementary videos S3 and S4, respectively. These videos show that the spin waves are emitted at the maxima of the oscillations in the skyrmion radius. This is remarkably similar to the experimental demonstration of spin-wave emission from an oscillating domain wall [44], where the oscillating wall ‘knocks’ the domain interior, stimulating spin waves [44]. It is clear from Fig.5(a,b) and Fig. 5(g,h) that the main difference between being driven at 57.4 GHz—where there is skyrmion motion—and at 14.3 GHz—where there is no skyrmion motion—is the amplitude of the emitted spin waves. It is well understood that the oscillations in radius at the out-of-phase breathing mode are considerably larger than at the in-phase breathing mode [39, 40]. We suggest that the primary factor determining the strength of the emitted spin waves is the magnitude of the change in radius. This is in turn affected by the magnitude of the driving field. However, for the same magnetic field amplitude, this will increase when tuning onto a skyrmion breathing mode resonant frequency, and will also be considerably larger for the out-phase breathing mode frequency.
To investigate this in more detail, we examine the amplitude of the variation in radius in the absence of an in-plane magnetic field. We do this because the deformation of the skyrmions—as shown in Fig. 1—will affect being able to accurately estimate the change in radius. Fig. 5(e,f) show snapshots of the same slice through the centre of the skyrmion at the same time as Fig. 5(c,d) at the same driving field frequency of 57.4 GHz, but without the in-plane field. Likewise, Fig. 5(k,l) show the same, but at a driving field frequency of 14.3 GHz, equivalent to Fig. 5(i,j) but without the in-plane magnetic field. Both of these sets of snapshots are supported by supplementary videos S5 and S6 which show the same data over the first 1 ns of the simulation. For a detailed description of all of the supplementary videos see supplementary note LABEL:supp:sec:videos.
In Fig. 5(m,n), we show a simultaneous study of the skyrmion velocity (Fig. 5(m)) and the magnitude of the variation in radius (Fig. 5(n)), both as a function of the magnitude of the driving microwave magnetic field at a frequency of 57.4 GHz. For details of how the average change in radius was calculated, see supplementary note LABEL:supp;sec:rad_var. In Fig. 5(m), at fields above $\sim$20 mT there is a clear linear relationship between the velocity and the driving field. However, below fields of $\sim$20 mT, the skyrmion velocity drops off faster towards 0 m/s, and there is no motion below driving fields of $\sim$3 mT. This region of no skyrmion motion is represented with the grey shaded box. Since this is a simulated system with no pinning, this is clearly an intrinsic effect of their being an insufficient spin-wave emission amplitude to drive motion. Likewise, examining the magnitude of the variation in radius in Fig. 5(n), we draw the same threshold field for skyrmion motion from Fig 5(m) also as a grey shaded box. For these magnetic parameters, we are then able to define a variation in radius of the skyrmions below which there is no motion of $\langle\Delta r_{\mathrm{Sk}}\rangle$ = 0.44 nm, or $\sim$3.5 % of the static skyrmion radius $r_{0}$.
To summarise, the motion of the skyrmions is primarily driven by the emission of spin waves. As we have shown, this spin wave emission occurs when the magnitude of the skyrmion expansion is at its widest. The results of Fig. 5 show that there is a minimum variation of skyrmion radius required to emit spin waves that are strong enough to drive motion. This radius variation will itself be dependent on the driving field; however, the relationship between radius variation and driving field will be different at the two resonant breathing modes. For this reason, we have determined empirically that it is more accurate to define a minimum variation in radius rather than a minimum driving field. This interpretation also explains the absence of skyrmion motion at the resonant frequency of the in-phase breathing mode in Fig. 4(a). The variation in skyrmion radius at this frequency is well understood to be significantly smaller for the same driving magnetic field amplitude than when driven at the out-of-phase resonant frequency [39, 40]. Performing the same analysis on the variation in radius of the in-phase breathing mode when driven by a out-of-plane excitation field of 20 mT gives a $\langle\Delta r_{\mathrm{Sk}}\rangle$ of 0.286$\pm$0.002 nm, below this threshold for motion.
IV.4 The effect of antiferromagnetic interlayer exchange coupling
Figure 6: (a) Skyrmion velocity as a function of frequency plotted for varying antiferromagnetic exchange coupling $\sigma_{\mathrm{RKKY}}$. Dashed lines show fits to an asymmetric Lorentzian function used to extract the peak velocity. All simulations were performed with static fields of 50 mT out-of-plane and 0.5 T in-plane, along with a microwave driving field of 20 mT. (b) The extracted peak velocities as a function of the coupling strength. The regime where at least one skyrmion is annihilated when driven on resonance is cross-hatched, while the coupling regime where the skyrmions move unstably is single hatched. The dashed line shows an inverse cube law as a guide to the eye for peak velocities in the high coupling strength regime.
With our new understanding of the coupling between skyrmion velocity and the frequency of the out-of-phase breathing mode, we now move to studying it as a function of the antiferromagnetic interlayer exchange coupling strength. The results of measuring the velocity as a function of frequency at different coupling strengths $\sigma_{\mathrm{RKKY}}$ are presented in Fig. 6(a). The data presented here has been trimmed around the frequency of the out-of-phase breathing mode for clarity, so the broad peak in velocity associated with the in-plane field is not shown. This however follows the same behaviour—albeit shifted in frequency—as the one for $\sigma_{\mathrm{RKKY}}=0.3\ \mathrm{mJ/m^{2}}$ shown in Fig. 4(a) and discussed in supplementary note LABEL:supp:sec:high_freq_mode. The extracted values of skyrmion velocity as a function of frequency are fitted to an asymmetric Lorentzian lineshape, chosen for its empirical match to the data, which allows us to extract a value for the peak velocity. This peak velocity as a function of $\sigma_{\mathrm{RKKY}}$ is then shown in Fig. 6(b).
In Ref. [43], Qiu et al. discuss the variation of the skyrmion velocity (driven by modulations in anisotropy) as a function of $\sigma_{\mathrm{RKKY}}$, and propose this as a way of modulating the skyrmion velocity. Our results show that while this variation in velocity with $\sigma_{\mathrm{RKKY}}$ is true at a fixed frequency, it is not fundamentally related to the value of $\sigma_{\mathrm{RKKY}}$. Instead, we have clearly shown that the magnitude of the velocity is correlated to the size of the oscillations in skyrmion size. This modulation of the skyrmion size is itself a result of tuning onto the intrinsic frequency of the out-of-phase breathing mode. Thus, by examining the peak velocity value for each $\sigma_{\mathrm{RKKY}}$ we can understand the true impact of the antiferromagnetic interlayer exchange coupling strength on the skyrmion motion. We define three principal regimes of the skyrmion velocity when driven on the resonant frequency.
First, we define the skyrmion annihilation regime for low values of $\sigma_{\mathrm{RKKY}}$ below $\sim 0.03\ \mathrm{mJ/m^{2}}$. Here the skyrmions are stable in the static case and skyrmion motion is still possible when driving with AC magnetic fields. However, when it is driven close to the resonant frequency, the oscillations of the skyrmions at this field strength are sufficiently large to drive the annihilation of the skyrmions in at least one of the layers. Clearly the strength of the antiferromagnetic coupling is insufficient to stabilise the two skyrmions as a topological object.
For intermediate values of $\sigma_{\mathrm{RKKY}}$ between $\sim 0.03\ \mathrm{mJ/m^{2}}$ and $<0.1\ \mathrm{mJ/m^{2}}$, the strength of the coupling is sufficient to prevent the annihilation of the skyrmions, however their oscillations are substantially more unstable than at higher coupling strengths, and the x-axis trajectories (the direction orthogonal to the principal direction of motion) of the two guiding centres drift apart by up to several nanometres over the course of the simulations. We suggest that this disparity between the two skyrmions restricts their ability to move together efficiently, and we see that the peak velocity over this narrow region of $\sigma_{\mathrm{RKKY}}$ is almost completely flat. In the simulation at $\sigma_{\mathrm{RKKY}}=0.08\ \mathrm{mJ/m^{2}}$ there is a small uptick in velocity as we transition to the high $\sigma_{\mathrm{RKKY}}$ regime.
For values of $\sigma_{\mathrm{RKKY}}$ from $0.1\ \mathrm{mJ/m^{2}}$ and up, we see a very clear inverse relationship between the peak skyrmion velocity and the strength of the antiferromagnetic interlayer exchange coupling. The two skyrmions move together with regular oscillations in size for the whole simulation space. The decrease in velocity with $\sigma_{\mathrm{RKKY}}$ can be clearly understood as a consequence of the constraining effect of the antiferromagnetic coupling strength, which as it increases will reduce the magnitude of the oscillations in skyrmion radius, and in turn reduce the velocity.
While experimental measurements of skyrmions and other magnetic textures are limited in synthetic antiferromagnets, several studies report observations of skyrmions and domain walls. In Table 1 we summarise the reported antiferromagnetic coupling strengths from a literature search of DMI stabilised textures in synthetic antiferromagnets. The fact that most of these reported values fall in the range of $\sigma_{\mathrm{RKKY}}$ in our study where we report skyrmion motion suggests that this method would be appropriate for driving SAF skyrmions in a future computing device.
Table 1: Compilation of reported antiferromagnetic interlayer exchange coupling strengths from several previous studies.
Material Magnetic Coupling Reference
System Texture Strength ($\mathrm{mJ/m^{2}}$)
Co/Ru/Pt/Co Skyrmions -0.23 [25]
Pt/Co/CoFeB/ Skyrmions -0.13 [32]
Ir/Co/CoFeB/W
Co/Ni/Co/Ru/ Domain Walls -0.5 [28]
Co/Ni/Co
[[[[CoFeB/Ru/Pt/ Domain Walls -0.018 [30]
[…]
V Electric Field Driven Skyrmion Motion
Figure 7: Examination of the effects of electric field on the skyrmion breathing and velocity at $\sigma_{\mathrm{RKKY}}=-0.3\ \mathrm{mJ/m^{2}}$. This is performed in an out-of-plane magnetic field of 50 mT, and in the case of the motion an in-plane magnetic field of 0.5 T. (a) The PSD of the whole system along with the two individual layers when excited with a modulation of the uniaxial anisotropy of $3\times 10^{4}\ \mathrm{J/m^{3}}$ is plotted on the left axis as a function of frequency. On the right vertical axis, the skyrmion velocity when driven by an electric field as a function of frequency is plotted. Panels (b) and (c) show the variation in radius of the two skyrmions when driven by a sinusoidal electric field at 14.3 GHz and 57.4 GHz, respectively.
Previous reports have studied skyrmion motion driven by a microwave electric field [43]. In this work they simulated a similar system with an in-plane magnetic field and drove the motion of skyrmions by adding a microwave frequency perturbation to the anisotropy energy term. This mirrors the voltage controlled magnetic anisotropy that has been demonstrated in thin film multilayers [48]. Ref. [43] reported skyrmion velocities as a function of frequency, demonstrating a peak in the skyrmion velocity. However, the relationship between velocity and the resonant frequency of the skyrmion breathing modes was not identified. The frequencies at which they observed their peak velocites were close to the resonant frequency of the in-phase breathing mode that we have shown in Fig. 2. To understand the differences between the mechanisms of magnetic-field driven skyrmion motion and the electric-field driven case previously reported [43], we finally perform some basic simulations using an electric field. In the same way as Ref. [43], the effects of an electric field are simulated by modulating the uniaxial anisotropy with an additional microwave frequency perturbation. We use the same value of $3\times 10^{4}\ \mathrm{J/m^{3}}$ for the magnitude of this modulation of the uniaxial anisotropy in all simulations. An interlayer exchange coupling of $\sigma_{\mathrm{RKKY}}=-0.3\ \mathrm{mJ/m^{2}}$ was used, well into the high $\sigma_{\mathrm{RKKY}}$ regime from Sec. IV.4. To match the results from Sec. IV.2, we perform these simulations in a static out-of plane field of 50 mT as well as an in-plane field of 500 mT in the case of the simulations of motion.
We first examine the breathing modes of the skyrmions, which is shown in Fig. 7(a) for both the individual layers of the system as well as for the system as a whole. Here the features of the PSD occur at the same frequencies as in the magnetic field driven-breathing case, however the presence or lack thereof of each mode in the different layers is different. There is a strong peak at 14.3 GHz in all three traces, which in the magnetic-field driven case originates from in-phase breathing of the two skyrmions. However, the presence of the peak in the response of the whole system is less straightforward to account for. Alternatively, there is a peak in the whole system at 57.4 GHz where there is only a faint trace of in the responses of the individual layers. This behaviour is intriguing and merits further investigation.
We also drive the motion of the skyrmions with an electric field of the same magnitude and varying frequency. The measured velocities as a function of the electric field frequency are also plotted on the right hand axis of Fig. 7(a). Here we observe a strong peak in velocity that is well coupled to the lower frequency breathing mode. Unlike the magnetic field driven case, there is no observable motion at the higher frequency mode, but there is some motion around the frequency of the mode excited by the in-plane field (see supplementary note LABEL:supp:sec:high_freq_mode). To understand this we present in Figs. 7(b) and (c) simulations of the radius variation of the skyrmions when driven by a sinusoidal electric field with no in-plane magnetic field at each of the principal frequencies from Fig. 7(a). At the lower 14.3 GHz peak, the radii of the two skyrmions varies massively with this electric field strength of $\langle\Delta r_{\mathrm{Sk_{1}}}\rangle=\langle\Delta r_{\mathrm{Sk_{2}}}\rangle=13.6\pm 0.2$ nm. Our hypothesis in Sec. IV.3 was that to stimulate skyrmion motion, a minimum variation in the skyrmion radius—of 0.44 nm for these micromagnetic energy terms—is required. The variation in radius at the in-phase resonant frequency is clearly large enough to drive motion. Likewise, this also explains the absence of motion at the other resonant breathing mode. The average variation in the radii of the two skyrmions when driven by an electric field at 57.4 GHz are $\langle\Delta r_{\mathrm{Sk_{1}}}\rangle=0.1768\pm 0.0001$ nm and $\langle\Delta r_{\mathrm{Sk_{2}}}\rangle=0.1821\pm 0.0001$ nm respectively, well below the threshold for motion.
VI Conclusion
To conclude, we have used micromagnetic simulations to demonstrate a new method for stimulating skyrmion motion in a synthetic antiferromagnet using only global magnetic fields. We have used a microwave frequency out-of-plane magnetic field to excite skyrmion breathing, which in the presence of a static in-plane field emits spin waves asymmetrically. Through characterisation of the breathing modes of the skyrmions in our systems we have understood the underlying relationship between skyrmion velocity and the frequency of the driving field, observing that the maximum skyrmion velocity occurs at the characteristic frequency of the out-of-phase skyrmion breathing mode.
We have further measured the variation in radius of the skyrmions at different driving fields to understand that there is a threshold magnitude of radius variation, only above which will skyrmion motion be driven.
The understanding of the coupling between the skyrmion velocity and the out-of-phase breathing mode has allowed us to properly understand the relationship between the velocity and the antiferromagnetic interlayer exchange coupling strength, where we define three different regimes of skyrmion motion depending on the magnitude of the antiferromagnetic exchange coupling.
Finally, we have compared our results to the case when driven by a microwave electric field. By doing this we have revealed both the true origin of the relationship between velocity and exchange coupling, and the reason for the lower frequency of skyrmion motion when driven by electric fields compared to magnetic fields.
These results will help to inform the design of future novel computing architectures based on the dynamics of skyrmions in synthetic antiferromagnets.
Acknowledgements.
This project was funded by the UK Government Department for Science, Innovation and Technology through National Measurement System funding (Metrology of complex systems for low energy computation). The project was also supported by EPSRC Grant No. EP/T006803/1.